The two numbers in ratio of 9:16 when on each number is increase by 15 the ratio become 2:3 what the ratio become if each of origins number decrease by 12
Answer by rothauserc(4718)

(9x+15)/(16x+15) = 2/3
:
3(9x+15) = 2(16x+15)
:
27x +45 = 32x +30
:
5x = 15
:
x = 3
:
9(3) = 27
:
16(3) = 48
:
27/48 is original fraction
:
27 -12 = 15
:
48 -12 = 36
:
*****************************
15/36 = 5/12 is the new ratio
*****************************
: